OpenStudy (anonymous):

How to solve for y, in a situation like: 1 = (constant - x)^2 + (constant - y)^2

OpenStudy (anonymous):

let a = constant #1, b=constant #2 expand the x term: 1=(a-x)^2 + (b-y)^2 1=(a^2-2ax+x^2) + (b-y^2) (b-y)^2=1-a^2+2ax-x^2 b-y = sqrt (1-a^2+2ax-x^2) y=b-sqrt (1-a^2+2ax-x^2)